Circular No. 8906

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

V574 PUPPIS D. K. Lynch, R. J. Rudy, T. R. Prater, A. M. Gilbert, and S. Mazuk,

The Aerospace Corporation; R. B. Perry, Langley Research Center, NASA; and R. C. Puetter, University of California, San Diego; report 0.47- to 2.5μm spectroscopy of V574 Pup (cf. IAUC 8443, 8445, 8643) on 2007 Dec. 15 UT using VNIRIS on the Lick 3-m telescope. The object is a super-soft x-ray source, and it continues in its coronal phase. Emission lines of [S IX] 1.252 μm, [Ca VIII] 2.3 μm, [Fe X] 637.5-nm, and [Fe XI] 789.2-nm were seen. He II 1.02 μm was stronger than He I 1.0830 μm. The [S III] and [O III] lines were weak.

COMET 8P/TUTTLE D. Schleicher, Lowell Observatory; and L. Woodney, California State

University, San Bernardino, obtained narrowband imaging of comet 8P on Dec. 15 and 16 using the Hall 1.1-m telescope at Lowell Observatory. Following the removal of a mean radial profile, a series of successive spiral arcs are seen in the sunward hemisphere for each of the brighter gas species (i.e., CN, C_{3}, and C_{2}). With over 8 hr of imaging per night, along with 4 hr on Dec. 14 (see IAUC 8903), they find a nucleus rotation period of ∼ 5.71 hr (estimated uncertainty 0.04 hr) based on 14 pairs of matching CN images from various rotational cycles. A single source appears to produce most, if not all, of the coma morphology.

COMET C/2007 T5 (GIBBS) Improved elliptical orbital elements for this comet (cf. IAUC 8880)

from MPEC 2007-Y07.

Epoch = 2008 May 14.0 TT

T = 2008 May 24.1765 TT  ω = 34.^{o}4057
e = 0.913398  Ω = 109.8429 }2000.0
q = 4.049340 AU  i = 45.6151

COMETS 194P/LINEAR AND 195P/HILL Comet P/2007 W2 = P/2000 B3 (cf. IAUC 8900) has been assigned the

permanent number 194P, and comet P/2006 W4 = P/1993 D1 (cf. IAUC 8779, 8902, 8903) has been assigned the permanent number 195P (cf. MPC 61390).

2007 December 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green